# Is my Data in your AI Model? Membership Inference Test with Application to Face Images

### v2

@@ IV-B3 MINT Models Training
Both architectures are chosen after extensive experimentation. It was found that the architectures did not need to be overly complex, as performance did not improve significantly with more complex architectures. In the experimental section (Section V-A), an experiment has been included to demonstrate the model’s performance versus the network complexity.
V Experiments and Results
This section presents the results achieved by our proposed MINT approach. As it involves binary classification between $\mathcal{D}$ and $\mathcal{E}$, and considers a balanced evaluation, the baseline accuracy (i.e., random guess) stands at 50%. Additionally, we remind that all results are derived from samples in $\mathcal{E}$ that do not belong to the databases used in training the MINT Model, as elaborated in Sec. IV-B.
The nomenclature used in this section is as follows::
  • FR Model ($M$ in Fig. 3): These are the Face Recognition models presented in Sec. IV-A. The number corresponds to the one specified in that section.
  • Model Outcome: Refers to the Output Embedding of the Face Recognition model in Sec. III-B.
  • Conv Layer X: Represents the AAD obtained from each of the four primary layers, as explained in Sec. IV-B2
  • Combination: Indicates the concatenation of the AAD obtained from the four primary layers.
@@ V Experiments and Results
V-A Results
Table I shows the classification accuracy results for the $3$ Face Recognition models using different information to train the Vanilla MINT Model $T$ (AAD and Model Outcomes). All of this is done under a scenario of high data available for training. In these results, we can observe a superiority in accuracy when combining all the AAD information (Combination). In terms of the AAD it is important to highlight that in FR Model 1, Conv Layer #2 yields the best results after Conv Layer #1, while in FR Model 2 and 3, Conv Layer #4 follows with better accuracy. This might be attributed to the fact that ResNet architectures are based on “Skip Connections”, allowing information to traverse through the network from first layers to the last ones [67].
TABLE I: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@ V-A Results
The results in this section pertain to the Baseline Case presented in Sec. IV-B, where the databases IJB-C, FDDB, and GANDiffFace are used in training the MINT Models, while Adience is used for evaluation. This is the case that provides a more robust evaluation (Sec. IV-B), and therefore the one mostly used throughout the experimental section. However, to demonstrate the robustness of the results, Table II showcases the same results as Table I with exchanged training and evaluation databases for the other two scenarios (see Sec. IV-B):
@@
As shown in Table II, the outcomes are slightly superior for Case A and slightly inferior for Case B compared to the Baseline Case, yet the the conclusions remain unchanged (results are presented as Case A / Case B).
TABLE II: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3). We display the results for Cases A and B (Case A/Case B in Table).
@@
In Table III, we can observe the classification accuracy when we vary the data available to train the Vanilla MINT model for the best obtained Auditable Data (combination of all convolutional layers). Naturally, accuracy decreases when we reduce the number of training samples, demonstrating that the patterns to be learned are not trivial. A noteworthy result is that in the $50K$ scenario, accuracy decreases very slightly with respect to $100K$. This result suggests that an intermediate number of training samples may be enough to achieve acceptable results. Another interesting result is what occurs in the $1K$ scenario. For FR Model 1, performance decreases by $18\%$. However, for FR Model 2 and FR Model 3, it decreases by only $8\%$ and $6\%$, respectively. From these results, we can infer that depending on the model $M$ (its training, architecture, loss function, etc.), it may be possible to develop our MINT Model $T$ with more or less training data.
TABLE III: Classification accuracy for the Vanilla MINT Model (Combination as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
In Table IV, we can observe the different classification accuracy results for the $3$ Face Recognition Models using the different information available to train the CNN MINT Model $T$ under the high data scenario. Here, the CNN is trained using the full activation blocks of each convolutional layer. In this case, just like before, the best results are achieved with Conv Layer #1 (i.e., the layer closest to the image domain), progressively decreasing as the layer gets closer to the end of the network (i.e., the layer closest to the output domain). It is important to note that, due to the different sizes of activation maps in the different layers ($H\times W$ depends on the depth at which you take the activation map), combining multiple layers to form the AAD in this CNN MINT would require a specific architecture or preprocessing, which is out of the scope of this paper. The Model Outcome is also not used, as explained in Sec. III-C, because, as we discussed there, it is vectorial information that does doesn’t fit in a CNN architecture in a straightforward way.
TABLE IV: Classification accuracy for various AAD using the CNN MINT model. MINT model trained with $100$K samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@
In Table V, we observe the impact of varying the number of available data used to train the CNN MINT Model $T$ for the best considered Auditable Data (Conv Layer #1). We see that for FR Model 1, a scenario with limited data for training the MINT Model negatively affect our results. However, for the other two FR Models, it has almost no effect on accuracy, achieving consistently high accuracy despite decreasing significantly the amount of training data. This indicates that for this MINT architecture depending on the model being audited, few samples ($500$ from $\mathcal{D}$) are enough to achieve good results. This result is very interesting as it implies that the auditor would require little cooperation from the developer.
TABLE V: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
To conclude we compare our proposed MINT model to the state of the art, in Table VI. As mentioned earlier, to the best of our knowledge, no state-of-the-art studies address the problem of detecting if given data was used for training a Face Recognition model. Therefore, making direct comparisons with other studies is not possible. Nevertheless, most studies around MIAs often use the classification output vector of the network [14, 28, 29, 16], or the output from the layers just before this output [30, 18]. These approaches are somewhat analogous to our Model Outcome (the Model Outcome of a Face Recognition Model is similar to the layers just before the output vector in a classification model), and thus, we use them as a reference for interpreting the state of the art, despite the environmental differences between the works around MIAs and our proposed MINT.
TABLE VI: Classification accuracy comparison between MIA and the two MINT approaches proposed in this work. We have adapted the MIA approach used in [18] using the Face Recognition Model outcome. Note that we used the original model instead of shadow models. ∗∗We used the Vanilla MINT model based on the combination of all convolutional layers. ∗∗∗We used the CNN MINT Model trained with the Conv Layer #1.
@@
Subsequently, we compare in Table VI the best published MIA results with the best results achieved in our present study. As observed, the methods proposed in the present study significantly outperform what could be considered the state of the art in Face Recognition MIAs, surpassing it for all three FR Models by more than $10\%$. Additionally, it is important to note that the results from the CNN MINT Model suggest even higher performance may be obtained if we combine information from all convolutional layers with specific architectures. Finally, Fig. 5 provides a comparison of the ROC curves for the three methods outlined in the table VI, along with the results for the Vanilla network with information from individual Conv Layers. An important leap is evident between MIA (purple curve) and our best result using the CNN MINT Model (black curve), which we have also observed in Table VI. One of the main criticisms discussed in [18] and [15] is that the evaluation protocol and metrics commonly used in MIA approaches are not adequate, and the results presented in previous studies suffer from a high False Positive Rate. In Fig. 5, we can see that this is not the case in our study.
As explained in Sec. IV-B3, an in-depth heuristic analysis (not fully reported here) has been conducted varying the complexity of the presented MINT Model architectures. From that study we can conclude that increasing the complexity of the architecture does not easily improve the performance. In our experiments, more complex architectures led to overfitting and a loss of performance in evaluation. To demonstrate this, Table VII presents results by varying the number of parameters of the MINT CNN architecture. We provide results with the original model as explained in Sec. IV-B3, another architecture with approximately three times the number of parameters, and the last one with ten times the number of parameters. We also include an experiment with a model three times simpler to demonstrate that the CNN requires enough complexity.
TABLE VII: Classification accuracy comparison while varying the complexity of the MINT CNN network used. CNN MINT architecture with ÷$\div 3$ parameters. ∗∗ Original MINT CNN architecture presented in IV-B3. ∗∗∗ CNN MINT architecture with ×$\times 3$ parameters. ∗∗∗∗ CNN MINT architecture with ×$\times 10$ parameters.
@@
VI Discussion
@@
It is important to contextualize the results achieved in the present study. While in Sec. V-A we provide a comparison with the state of the art in MIAs, this is for a simple interpretation as there is no specific work in the state of the art that aims at detecting if given data was used for training on the task of Face Recognition. In the state of the art of MIAs, results are available for classification tasks with very simple databases such as CIFAR-10 [68], where accuracy reaches up to 0.631 [69]. In the present study, we present results for the Face Recognition task. The complexity of this task is much higher, and with more relevant real-world application. Furthermore, the results achieved with the proposed MINT approach reach up to $90\%$, significantly surpassing those for CIFAR-10.
@@
On the other hand, it is important to emphasize the context of our study and the results achieved. As explained in Sec. III, existing MIAs and our proposed MINT operate under different environmental conditions. These differences arise from the distinct application scenarios of both technologies. While MIA is designed as an attack, MINT is conceived as an auditing tool motivated on the recent legislation and aims for user transparency. As a result, MIAs require the creation of shadow models to simulate the behavior of the original model to attack, whereas in MINT we have direct access to the original model. This leads to divergent outcomes and the potential utilization of information in MINT that is not particularly useful in MIA. For instance, the use of intermediate activations seems not to yield substantial improvements in MIAs [18, 31, 30]. Nevertheless, as demonstrated in our study, appropriately leveraging this information significantly enhances performance in MINT. This implies that ineffective methods in MIA do not necessarily perform poorly in MINT, suggesting opportunities for technology-specific advancements. As an example, exploring the use of gradients could be a promising avenue for future research in MINT.
VII Conclusions
@@
This work opens numerous avenues for future research. A direct path is to enhance the presented work by employing a different architecture for the MINT Model, such as using Visual Transformers or incorporating additional information like the gradients of the data. The study demonstrates the effectiveness of the MINT Model even with only $1$K training samples; a subsequent step would be to explore unsupervised training and identify patterns in the activations in an unsupervised manner, eliminating the need for developer involvement. Another future direction could involve training the MINT Model alongside the Audited Model with a joint loss function to optimize performance for both tasks simultaneously. This would necessitate active collaboration from the model developer, leading to what we might call Active Membership Inference Testing, as opposed to the Passive Membership Inference Testing presented in this work. Additional lines of research include applying this technology to other applications, such as image generation models [70, 71], or to different types of data, such as text [72], and studying MINT on widely available LLMs [73].

### v3

@@ 4.2.3 MINT Models Training
Both architectures are chosen after extensive experimentation. It was found that the architectures did not need to be overly complex, as performance did not improve significantly with more complex architectures. In the experimental section (Section 5.1), an experiment has been included to demonstrate the model’s performance versus the network complexity.
5 Experiments and Results
This section presents the results achieved by our proposed MINT approach. As it involves binary classification between $\mathcal{D}$ and $\mathcal{E}$, and considers a balanced evaluation, the baseline accuracy (i.e., random guess) stands at 50%. Additionally, we remind that all results are derived from samples in $\mathcal{E}$ that do not belong to the $\mathcal{E}$ databases used in training the MINT Model, as elaborated in Sec. 4.2.
The nomenclature used in this section is as follows:
  • 1.
    FR Model ($M$ in Fig. 3): These are the Face Recognition models presented in Sec. 4.1. The number corresponds to the one specified in that section.
  • 2.
    Model Outcome: Refers to the Output Embedding of the Face Recognition model in Sec. 3.2.
  • 3.
    Conv Layer X: Represents the AAD obtained from each of the four primary layers, as explained in Sec. 4.2.2
  • 4.
    Combination: Indicates the results using the AAD obtained from the four primary layers. The combination, including the Model Outcome as well, yields exactly the same results.
  • 5.
@@ 5 Experiments and Results
5.1 Results
Table 1 shows the classification accuracy results for the $3$ Face Recognition models using different information to train the Vanilla MINT Model $T$ (AAD and Model Outcomes). All of this is done under a scenario of high data available for training. In these results, we can observe a superiority in accuracy when combining all the AAD information (Combination). In terms of the AAD it is important to highlight that in FR Model 1, Conv Layer #2 yields the best results after Conv Layer #1, while in FR Model 2 and 3, Conv Layer #4 follows with better accuracy. This might be attributed to the fact that ResNet architectures are based on “Skip Connections”, allowing information to traverse through the network from first layers to the last ones [76].
Table 1: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@ 5.1 Results
The results in this section pertain to the Baseline Case presented in Sec. 4.2, where the databases IJB-C, FDDB, and GANDiffFace are used in training the MINT Models, while Adience is used for evaluation. This is the case that provides a more robust evaluation (Sec. 4.2), and therefore the one mostly used throughout the experimental section. However, to demonstrate the robustness of the results, Table 2 showcases the same results as Table 1 with exchanged training and evaluation databases for the other two scenarios (see Sec. 4.2):
  • 1.
@@
As shown in Table 2, the outcomes are slightly superior for Case A and slightly inferior for Case B compared to the Baseline Case, yet the the conclusions remain unchanged (results are presented as Case A / Case B).
Table 2: Classification accuracy for various AAD and Model Outcomes using the Vanilla MINT Model. MINT model trained with $100K$ samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3). We display the results for Cases A and B (Case A/Case B in Table).
@@
In Table 3, we can observe the classification accuracy when we vary the data available to train the Vanilla MINT model for the best obtained Auditable Data (combination of all convolutional layers). Naturally, accuracy decreases when we reduce the number of training samples, demonstrating that the patterns to be learned are not trivial. A noteworthy result is that in the $50K$ scenario, accuracy decreases very slightly with respect to $100K$. This result suggests that an intermediate number of training samples may be enough to achieve acceptable results. Another interesting result is what occurs in the $1K$ scenario. For FR Model 1, performance decreases by $18\%$. However, for FR Model 2 and FR Model 3, it decreases by only $8\%$ and $6\%$, respectively. From these results, we can infer that depending on the model $M$ (its training, architecture, loss function, etc.), it may be possible to develop our MINT Model $T$ with more or less training data.
Table 3: Classification accuracy for the Vanilla MINT Model (Combination as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
In Table 4, we can observe the different classification accuracy results for the $3$ Face Recognition Models using the different information available to train the CNN MINT Model $T$ under the high data scenario. Here, the CNN is trained using the full activation blocks of each convolutional layer. Once again, the best performance is achieved with the Combination of all Conv Layers. Similarly, as observed previously, the next best results are achieved with Conv Layer #1 (i.e., the layer closest to the image domain), progressively decreasing as the layer gets closer to the end of the network (i.e., the layer closest to the output domain). The Model Outcome is not used, as explained in Sec. 3.3, because, as we discussed there, it is vectorial information that doesn’t fit in a CNN architecture in a straightforward way. For these CNN MINT Models, training with the combination of all the AADs significantly increases the computational complexity (approximately fourfold), while yielding only a modest improvement of around 1% in accuracy. Consequently, for the remaining experiments, we proceed with the CNN MINT Model trained using Conv Layer 1 as the AAD.
Table 4: Classification accuracy for various AAD using the CNN MINT model. MINT model trained with $100$K samples ($1\%$ of the total Face Recognition Model training set for FR Model 1 and $0.3\%$ for FR Models 2 and 3).
@@
Combination 0.88 0.90 0.90
In Table 5, we observe the impact of varying the number of available data used to train the CNN MINT Model $T$ for the considered Auditable Data (Conv Layer #1). We see that for FR Model 1, a scenario with limited data for training the MINT Model negatively affect our results. However, for the other two FR Models, it has almost no effect on accuracy, achieving consistently high accuracy despite decreasing significantly the amount of training data. This indicates that for this MINT architecture depending on the model being audited, few samples ($500$ from $\mathcal{D}$) are enough to achieve good results. This result is very interesting as it implies that the auditor would require little cooperation from the developer. A similar situation occurs in Table 3 with the Vanilla MINT Model. The reason behind this is that, in Model 1, the activation patterns indicating whether a sample was used for training or not are more complex than in Models 2 and 3. As a result, the MINT Model requires more training data compared to the other two models to perform effectively.
Table 5: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) in the various data provided scenarios discussed for the three Face Recognition (FR) models.
@@
To conclude we compare our proposed MINT model to the state of the art, in Table 6. As mentioned earlier, to the best of our knowledge, there are no studies addressing the problem of detecting whether specific data was used for training state-of-the-art Face Recognition models. Therefore, making direct comparisons with other studies is not possible. Nevertheless, most studies around MIAs often use the classification output vector of the network [15, 38, 39, 17], or the output from the layers just before this output [40, 19]. These approaches are somewhat analogous to our Model Outcome (the Model Outcome of a Face Recognition Model is similar to the layers just before the output vector in a classification model), and thus, we use them as a reference for interpreting the state of the art, despite the environmental differences between the works around MIAs and our proposed MINT.
Table 6: Classification accuracy comparison between MIA and the two MINT approaches proposed in this work. We have adapted the MIA approach used in [19] using the Face Recognition Model outcome. Note that we used the original model instead of shadow models. ∗∗We used the Vanilla MINT model based on the combination of all convolutional layers. ∗∗∗We used the CNN MINT Model trained with the Conv Layer #1.
@@
Subsequently, we compare in Table 6 the best published MIA results with the best results achieved in our present study. As observed, the methods proposed in the present study significantly outperform what could be considered the state of the art in Face Recognition MIAs, surpassing it for all three FR Models by more than $10\%$. Finally, Fig. 5 provides a comparison of the ROC curves for the three methods outlined in the table 6, along with the results for the Vanilla network with information from individual Conv Layers. An important leap is evident between MIA (purple curve) and our best result using the CNN MINT Model (black curve), which we have also observed in Table 6. One of the main criticisms discussed in [19] and [16] is that the evaluation protocol and metrics commonly used in MIA approaches are not adequate, and the results presented in previous studies suffer from a high False Positive Rate. In Fig. 5, we can see that this is not the case in our study.
As explained in Sec. 4.2.3, an in-depth heuristic analysis (not fully reported here) has been conducted varying the complexity of the presented MINT Model architectures. From that study we can conclude that increasing the complexity of the architecture does not easily improve the performance. In our experiments, more complex architectures led to overfitting and a loss of performance in evaluation. To demonstrate this, Table 7 presents results by varying the number of parameters of the MINT CNN architecture. We provide results with the original model as explained in Sec. 4.2.3, another architecture with approximately three times the number of parameters, and the last one with ten times the number of parameters. We also include an experiment with a model three times simpler to demonstrate that the CNN requires enough complexity. Finally, we also present results using two additional architectures widely employed in the state of the art in computer vision. Specifically, the results obtained using a ResNet50 [76] and a ViT [77] as MINT Models.
Table 7: Classification accuracy comparison while varying the complexity of the MINT CNN network used. CNN MINT architecture with ÷$\div 3$ parameters. ∗∗ Original MINT CNN architecture presented in 4.2.3. ∗∗∗ CNN MINT architecture with ×$\times 3$ parameters. ∗∗∗∗ CNN MINT architecture with ×$\times 10$ parameters. ∗∗∗∗∗ ResNet50 [76] as CNN MINT architecture. ∗∗∗∗∗∗ ViT [77] as CNN MINT architecture
@@
MINT ResNet50 $25.6$M 0.86 0.89 0.89
MINT ViT $86$M 0.86 0.89 0.90
6 Why Face Recognition? Exploring its Relevance and Broader Applications
In this initial work, the decision was made to focus on the field of computer vision (CV) due to the wide availability of datasets and models and the significant relevance of CV within AI. However, applying this technology to fields like NLP is also highly interesting, as NLP often requires an immense amount of data, which can sometimes be subject to copyright protection.
Within computer vision, face recognition (FR) is one of the areas where leveraging massive amounts of data could be particularly beneficial for improving systems [78, 79]. This need may tempt developers to use data without proper authorization. FR is also highly sensitive to variability and changes, often necessitating model retraining to adapt to new situations—e.g., the use of masks during the COVID-19 pandemic, which prompted adjustments in face recognition systems [78]. The need for rapid updates can further encourage the use of data without adequate consent. Additionally, facial images in CV are highly sensitive, often met with public opposition [80, 21, 22, 23] and protected by various legal frameworks [24, 25, 26]. These factors collectively motivated us to introduce MINT in the FR domain.
Nevertheless, within CV, we have also applied MINT to image classification tasks. Although image classification might be less visually appealing than face recognition and its utility may seem less obvious, it remains critical. Many images, not limited to faces, may belong to individuals and be protected under copyright—e.g., landscape photos taken by a photographer—making their unauthorized use a potential violation of legislation and individual rights. Moreover, MINT is a valuable tool for international competitions, where the goal is often to improve model classification performance, and it is crucial that MINT performs effectively in these tasks to identify whether participants have trained models with unauthorized data. Experiments in the domain of image classification can be found in Appendix A.
7 Real-World Deployment of MINT: Challenges and Legal Implications
7.1 Obfuscation Methods
MINT is a tool designed to identify data that has been used to train a model, aiming to protect citizen rights and promote safer and more trustworthy AI. It seeks to prevent developers from leveraging data without the required permissions. However, this raises an important question: Can developers use unauthorized data and deliberately conceal its use from MINT Models?
In this work, we demonstrate that detection of training data is feasible even with state-of-the-art models. Nevertheless, developers aware of MINT’s capabilities could design new training techniques to obscure the presence of unauthorized data, which would necessitate further advancements in MINT methods. This interplay opens a future research avenue, requiring MINT approaches to evolve alongside emerging concealment techniques.
7.2 Deployment and Legal Implications
The deployment process of MINT could work as follows: a developer releases a new model $M$ (e.g. a FR Model), and an auditing entity (e.g., the EU or a private organization) trains the MINT model $T$ and makes it publicly available. This auditing entity could verify whether certain popular databases or internet-scraped data were used in training the model $M$. Simultaneously, users can test the model to determine if their images, such as those shared on social media, were used without consent and report such findings to the auditing entity. If it is detected that a developer trained a model using data without proper authorization, severe legal, ethical, and reputational consequences could follow.
From a legal perspective, particularly regarding biometric data like facial images, many countries enforce regulations that protect these sensitive data types, such as the GDPR [24], CCPA [25], and BIPA [26]. These laws can impose substantial financial penalties in addition to potential class-action lawsuits. Furthermore, if the data were sourced from databases or platforms, this could violate their terms of service and copyright laws. Developers may also face individual or collective lawsuits from affected individuals.
On the other hand, developers would also face ethical and reputational consequences, leading to a loss of trust from users in both the developers and their associated companies, as well as generating negative publicity.
Furthermore, legal penalties could include model withdrawal or access restrictions, while reputational harm might make it difficult for developers to launch future models that users trust and adopt.
8 Discussion
@@
It is important to contextualize the results achieved in the present study. While in Sec. 5.1 we provide a comparison with the state of the art in MIAs, this is for a simple interpretation as there is no specific work that aims at detecting if given data was used for training on state-of-the-art Face Recognition Models. In the state of the art of MIAs, results are available for classification tasks with small databases such as CIFAR-10 [81] where accuracy reaches up to 0.631 [82] or Yale Face [83] where accuracy is 0.558 [28]. The complexity of state-of-the-art Face Recognition models and databases is much higher, and with more relevant real-world application. Furthermore, the results achieved with the proposed MINT approach reach up to $90\%$.
@@
On the other hand, it is important to emphasize the context of our study and the results achieved. As explained in Sec. 3, existing MIAs and our proposed MINT operate under different environmental conditions. These differences arise from the distinct application scenarios of both technologies. While MIA is designed as an attack, MINT is conceived as an auditing tool motivated on the recent legislation and aims for user transparency. As a result, MIAs require the creation of shadow models to simulate the behavior of the original model to attack, whereas in MINT we have direct access to the original model. This leads to divergent outcomes and the potential utilization of information in MINT that is not particularly useful in MIA. For instance, the use of intermediate activations seems not to yield substantial improvements in MIAs [19, 41, 40]. Nevertheless, as demonstrated in our study, appropriately leveraging this information significantly enhances performance in MINT. This implies that ineffective methods in MIA do not necessarily perform poorly in MINT, suggesting opportunities for technology-specific advancements. As an example, exploring the use of gradients could be a promising avenue for future research in MINT.
9 Conclusions
@@
This work opens numerous avenues for future research. A direct path is to enhance the presented work by incorporating additional information like the gradients of the data. The study demonstrates the effectiveness of the MINT Model even with only $1$K training samples; a subsequent step would be to explore unsupervised training and identify patterns in the activations in an unsupervised manner, eliminating the need for developer involvement. Another future direction could involve training the MINT Model alongside the Audited Model with a joint loss function to optimize performance for both tasks simultaneously. This would necessitate active collaboration from the model developer, leading to what we might call Active Membership Inference Testing, as opposed to the Passive Membership Inference Testing presented in this work. Additional lines of research include applying this technology to other applications, such as image generation models [84, 85], or to different types of data, such as text [86], and studying MINT on widely available LLMs [87].